student174:

Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar. A rectangle is 12 feet long and 5 feet wide. If the length of the rectangle is increased by 25% and the width is decreased by 20%, what is the change in the area of the rectangle? The change in the area of the rectangle is

Convert:

Length: 12 (0.25) =3 12+3=15 Width: 5 (0.2)=1 5-1=4 Area from original: 12*5=60 Area from new one: 15*4=60 Answer: No change in area